the store sells cashews for 6.00 per pound and peanuts for 3.00 per pound the manager decides to mix 60 pounds with some cashews and sell the mixture for 4.oo per pound how many pounds should be mixed with the cashews to produce the same revenue. I keep getting this wrong.
Answer by scott8148(6628)

3(x) + 6(60 - x) = 4(60)

3x + 360 - 6x = 240

120 = 3x ___ 40 = x